OpenStudy (anonymous):

Can someone help me with this problem Emily and Greta have volunteered to fold and staple a box of flyers in 3 hours. Greta works more efficently and can get the job done in 2 hours. How long will it take the girls to complete the job if they work togther?

OpenStudy (anonymous):

6/5

OpenStudy (anonymous):

Uhm hm1618 wat is 6/5?

OpenStudy (anonymous):

1 hour 12 mins

OpenStudy (anonymous):

thats what i got 1. ut im not sure.

OpenStudy (anonymous):

no I think I was right

OpenStudy (anonymous):

They want me to give the answer in minutes

OpenStudy (anonymous):

\[\frac{x}{3}+\frac{x}{2}=1\]\[3x+2x=6\]\[5x=6\]\[x=\frac{6}{5}\]

OpenStudy (anonymous):

6/5 is a improper fraction, correct?

OpenStudy (anonymous):

\[1\frac{1}{5}\] = 1 hours , 12 minutes

OpenStudy (radar):

I don't think there is enough facts presented here. 1. The actual time for both girls working together is not given. 2. The time to complete the job is not given 3 The time Emily takes is not given except we know that she would take longer working alone than Greta, in other words greater than two hours. 4. It is not stated how more efficient Greta is than Emily. 5. The key would be how long would it take Emily on her own. If Emily took 4 hr alone then her and Greta could get it done in 80 minutes, if it took her 2.5 hrs, then it could be done 1.111 hr. I don't see where the information this given The fact they have volunteered to do it in three hours and Greta can accomplish the job in 2 hours is just clutter.